OpenStudy (anonymous):

find the domain: f(x)=x^2+1

OpenStudy (angela210793):

D=b^2-4ac=-4ac=-4*1*1=-4

OpenStudy (amistre64):

where did you last see it?

OpenStudy (angela210793):

hahaha :D :D...

OpenStudy (anonymous):

I have 8 of these that need answers....

OpenStudy (amistre64):

the domain is all valid inputs; this function has no restrictions so all real numbers

OpenStudy (angela210793):

U have the formula now....so i guess u can do them by urself :) At least try 2st :)

OpenStudy (angela210793):

1st*

OpenStudy (anonymous):

Thank you, I will

OpenStudy (amistre64):

teh discriminate doesnt define a domain tho

OpenStudy (amistre64):

discriminate tells you if it hits the x axis; but domain tells you what you can use to find y

OpenStudy (amistre64):

a graph that floats above the x axis still has a domain of all valid x's; otherwise you simply couldnt have the graph

OpenStudy (anonymous):

The domain is R

OpenStudy (amistre64):

(-inf,inf)

OpenStudy (anonymous):

R -> Real no.s

OpenStudy (amistre64):

you gotta worry about even root, logs and fractions....

OpenStudy (amistre64):

and some tan,cot, inverse trig functions

OpenStudy (amistre64):

and spiders; if you come across a spider, you got issues lol

OpenStudy (anonymous):

(brain smoking)

OpenStudy (anonymous):

@amistre64: Don't confuse anyone! :P

OpenStudy (amistre64):

but... its what I do best ;)

OpenStudy (angela210793):

@amistre64 Agree :P

OpenStudy (anonymous):

lol

OpenStudy (angela210793):

@amistre64 May i ask u smth? (If u don't mind?) :)

OpenStudy (amistre64):

sure ....

OpenStudy (angela210793):

Thnx ^_^ (it's not abt math) R u indian?

OpenStudy (amistre64):

..... no. Mistretta is more Sicilian in origin

OpenStudy (anonymous):

Are you Indian?

OpenStudy (amistre64):

im a native floridain tho... does that count :)

OpenStudy (angela210793):

Oh..Italy....i knew it u weren't from America...(europeans r smarter) No i'm not

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

I'm Indian?y?

OpenStudy (amistre64):

usnret; the answer is... im not indian lol

OpenStudy (anonymous):

lol

OpenStudy (angela210793):

Nothing...i was just wondering if he was indian..cause indians are grrrrrrreat in math....

OpenStudy (angela210793):

hahaha :D

OpenStudy (amistre64):

The domain of your function tells you all the valid inputs that make it work right. Those values for this function have no restrictions so any number will work

OpenStudy (anonymous):

wish i was indian..lol

OpenStudy (angela210793):

haha me 2 :) :)

OpenStudy (anonymous):

I feel so proud! :P

OpenStudy (amistre64):

us mafia guys knows how to count ...

OpenStudy (anonymous):

what murders?

OpenStudy (angela210793):

Is it Cosa nostra in Sicily or Camorra?

OpenStudy (amistre64):

nownow...not just counting kills; we also do money laundering, counterfeiting, and extortion

OpenStudy (amistre64):

hey vinney, why you only give me 32%, you think im stupid? boom boom, now i gots 100%

OpenStudy (anonymous):

I don't understand your language, I'm just 16, no offense please,underage.

OpenStudy (anonymous):

:P

OpenStudy (angela210793):

I don't understand him 2 O.o

OpenStudy (amistre64):

hey, i only use words I find in the dictionary lol

OpenStudy (anonymous):

But where the hell do you find such 'dictionary' dats the question!

OpenStudy (angela210793):

how come google doesn't translate it then? O.o

OpenStudy (amistre64):

lol... casue i cant type lol

OpenStudy (amistre64):

gimme a button and a stick and I can beat any video game; trying to keep track of 104 keys? lol aint gonna happen

OpenStudy (anonymous):

Then you may try typign slwoyl!

OpenStudy (amistre64):

oh i type slowly, its just that these fat fingers and this tiny little keyboard just dont agree to distance

OpenStudy (anonymous):

Hey bye gn, its 1:00 AM ova here! toodles

OpenStudy (angela210793):

I'm going too,will watch CSI-NY :) Good night Amogh :) and have a good evening Amistre :)

OpenStudy (amistre64):

Ciao yall ;) im stickin round for another hour waiting for stats class